Circular No. 5766

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA OPHIUCHI 1993 R. H. McNaught, Anglo-Australian Observatory, provides the following

position for N Oph 1993 from a film taken with the Uppsala Southern Schmidt telescope on Apr. 17.5 UT: α = 17^{h}22^{m}04^{s}.41, δ = −23^{o}08′32′′.2 (equinox 1950.0; uncertainty 0′′.3 in each coordinate). This position does not correspond to any object on the SERC J survey plate (limiting mag ∼ 21.5) taken 1976 May 29, on a UKST I plate taken 1980 Aug. 5, on an ESO B plate taken 1976 Apr. 5, or on an ESO R plate taken 1984 July 25. The nearest star measured from the J survey was at mag ∼ 18.5 located 2′′.6 east and 3′′.2 south of the nova. A star of mag ∼ 12 is located at α = 17^{h}21^{m}58^{s}.22, δ = −23^{o}07′42′′.9.

Photometry by A. C. Gilmore, Mt. John University Observatory, obtained in poor conditions with the 0.6-m f/13 reflector (reference star HR 6473): Apr. 18.49 UT, V = 9.66, B–V = +0.79, U–B = –0.17.

SUPERNOVA 1993J IN NGC 3031 H. U. Zimmermann, Max-Planck-Institut für Extraterrestrische Physik

(MPE); W. Lewin and E. Magnier, Massachusetts Institute of Technology; P. Predehl, G. Hasinger, W. Pietsch, B. Aschenbach, and J. Trumper, MPE; G. Fabbiano, Center for Astrophysics; J. van Paradijs, University of Amsterdam; L. Lubin, Princeton University; and R. Petre, Goddard Space Flight Center, communicate: “SN 1993J has been observed with the ROSAT Observatory on six occasions between Apr. 3 (IAUC 5748 and 5750) and Apr. 13. The exposures were made on Apr. 3, 7, 9, 10, 12, and 13 and lasted between 1000 and 3000 s each. A linear fit to the count rates (0.1-2.4 keV) shows a decrease of ∼ 30 percent in 10 days. Our observations are inconsistent with a constant count rate at a confidence level higher than 99.9 percent. This seems to support models that predict a decay in the x-ray intensity on timescales of a few weeks to months.”

COMET SHOEMAKER-LEVY (1993e) Total visual magnitude estimates (cf. IAUC 5745): Apr. 1.03 UT, 13.9

(H. Mikuž, Ljubljana, Slovenia, 0.20-m Baker-Schmidt camera + CCD + V filter); Apr. 1.73, 14.0 (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector); 12.28, 13.3 (A. Hale, Las Cruces, NM, 0.41-m reflector); 14.13, 12.8 (J. E. Bortle, Stormville, NY, 0.50-m reflector); 14.59, 13.9 (Nakamura).

1993 April 18 Daniel W. E. Green